OpenStudy (anonymous):

Solve the system by substitution. x + 3y = –4 2x + 2y = –8

OpenStudy (luigi0210):

Solve for x and then plug that into the second equation

OpenStudy (anonymous):

A ) x + 3y = -4 B) 2x + 2y = -8 A + B = ? then you can solve it...

OpenStudy (anonymous):

3x+5y+-12

OpenStudy (luigi0210):

@PFEH.1999 it says by substitution :P

OpenStudy (anonymous):

Oh sorry!I didn't read the question! SORRY!

OpenStudy (anonymous):

\[x + 3y = –4 \implies x= -3y-4 \implies 2(-3x-4)+2y=-8 \implies y=0\]\[\implies x=-4 \implies (x,y)=(-4,0)\]

OpenStudy (anonymous):

x=-3y-4 2(-3y-4)+2y=8 Решите для Y

OpenStudy (anonymous):

I have an error in the first line, should be -3y, not -3x.